Seventy-five percent of adult smokers started smoking before turning 18 years old. If 20 adult smokers are randomly chosen, find the probability that at most 6 of them started smoking after they turned 18.
Answer #1

p = 0.75

n = 20

P(X = x) = 20Cx * 0.75x * (1 - 0.75)20-x

P(X < 6) = P(X = 0) + P(X = 1) + P(X = 2) + P(X = 3) + P(X = 4) + P(X = 5) + P(X = 6)

= 20C0 * 0.750 * 0.3520 + 20C1 * 0.751 * 0.2519 + 20C2 * 0.752 * 0.2518 + 20C3 * 0.753 * 0.2517 + 20C4 * 0.754 * 0.2516 + 20C5 * 0.755 * 0.2515 + 20C6 * 0.756 * 0.2514

= 0.000030